$SPAD/src/input zimmer.input

Timothy Daly December 3, 2016

Abstract

These examples come from Postel and Zimmermann’s paper in the 5th Rhine Conference on Computer Algebra.

Contents

2

— * —

)set break resume
)sys rm -f zimmer.output
)spool zimmer.output
)set message test on
)set message auto off
-- First Order Equations
--S 1 of 177
u := operator 'u
--R
--R
--R   (1)  u
--R                                                          Type: BasicOperator
--E 1
--S 2 of 177
ode := (x^4-x^3)*D(u x,x) + 2*x^4*u(x) = x^3/3 + C
--R
--R
--R                                   3
--R          4    3  ,        4      x  + 3C
--R   (2)  (x  - x )u (x) + 2x u(x)= -------
--R                                     3
--R                                          Type: Equation(Expression(Integer))
--E 2
--S 3 of 177
solve(ode,u,x)
--R
--R
--R                         3     2                     - 2x
--R                       2x  - 3x  + 6C              %e
--R   (3)  [particular= ------------------,basis= [-----------]]
--R                        4      3      2          2
--R                     12x  - 24x  + 12x          x  - 2x + 1
--IType: Union(Record(particular: Expression(Integer),...
--E 3
)clear all

--S 4 of 177
u := operator 'u
--R
--R
--R   (1)  u

--R                                                          Type: BasicOperator
--E 4
--S 5 of 177
ode := -D(u x,x)/2 + u(x) = sin(x)
--R
--R
--R           ,
--R        - u (x) + 2u(x)
--R
--R   (2)  ---------------= sin(x)
--R               2
--R                                          Type: Equation(Expression(Integer))
--E 5
--S 6 of 177
solve(ode,u,x)
--R
--R
--R                     4sin(x) + 2cos(x)           2x
--R   (3)  [particular= -----------------,basis= [%e  ]]
--R                             5
--IType: Union(Record(particular: Expression(Integer),...
--E 6
)clear all

--S 7 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 7
--S 8 of 177
ode := D(y x,x) = y(x)/(y(x)*log(y(x)) + x)
--R
--R
--R         ,            y(x)
--R   (2)  y (x)= -----------------
--R               y(x)log(y(x)) + x
--R                                          Type: Equation(Expression(Integer))
--E 8
--S 9 of 177
solve(ode,y,x)
--R
--R

--R                     2
--R        y(x)log(y(x))  - 2x
--R   (3)  -------------------
--R               2y(x)
--R                                         Type: Union(Expression(Integer),...)
--E 9
)clear all

--S 10 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 10
--S 11 of 177
ode := 2*y(x)*D(y x,x)^2 -2*x*D(y x,x)-y(x) = 0
--R
--R
--R              ,   2      ,
--R   (2)  2y(x)y (x)  - 2xy (x) - y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 11
--S 12 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 12
)clear all

--S 13 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 13

--S 14 of 177
ode := D(y x,x) + y(x) = y(x)^3*sin(x)
--R
--R
--R         ,                3
--R   (2)  y (x) + y(x)= y(x) sin(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 14
--S 15 of 177
solve(ode,y,x)
--R
--R
--R               2              2
--R        - 4y(x) sin(x) - 2y(x) cos(x) + 5
--R   (3)  ---------------------------------
--R                         2  2x
--R                    5y(x) %e
--R                                         Type: Union(Expression(Integer),...)
--E 15
)clear all

--S 16 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 16
--S 17 of 177
P := operator 'P
--R
--R
--R   (2)  P
--R                                                          Type: BasicOperator
--E 17
--S 18 of 177
Q := operator 'Q
--R
--R
--R   (3)  Q
--R                                                          Type: BasicOperator
--E 18

--S 19 of 177
ode := D(y x,x) + P(x)*y(x) = Q(x)*y(x)^n
--R
--R
--R         ,                        n
--R   (4)  y (x) + P(x)y(x)= Q(x)y(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 19
--S 20 of 177
solve(ode,y,x)
--R
--R
--R   (5)  "failed"
--R                                                    Type: Union("failed",...)
--E 20
--S 21 of 177
solve(eval(ode,n=1),y,x)
--R
--R
--R                                    x
--R                                  ++
--I                                  |   (Q(%J) - P(%J))d%J
--R                                 ++
--R   (6)  [particular= 0,basis= [%e                        ]]
--IType: Union(Record(particular: Expression(Integer),...
--E 21
--S 22 of 177
solve(eval(ode,n=2),y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   Function not supported by Risch d.e.
--R
--R   Continuing to read the file...
--R
--E 22
--S 23 of 177
solve(eval(ode,n=%pi),y,x)
--R
--R
--R   (7)  "failed"
--R                                                    Type: Union("failed",...)
--E 23

--S 24 of 177
solve(eval(ode,n=%e),y,x)
--R
--R
--R   (8)  "failed"
--R                                                    Type: Union("failed",...)
--E 24
--S 25 of 177
solve(eval(ode,n=sqrt(2)),y,x)
--R
--R
--R   (9)  "failed"
--R                                                    Type: Union("failed",...)
--E 25
)clear all

--S 26 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 26
--S 27 of 177
ode := (x^2-1)*D(y x,x)^2 - 2*x*y(x)*D(y x,x)+(y x)^2 - 1 = 0
--R
--R
--R          2      ,   2           ,          2
--R   (2)  (x  - 1)y (x)  - 2x y(x)y (x) + y(x)  - 1= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 27
--S 28 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 28
)clear all

--S 29 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 29
--S 30 of 177
f := operator 'f
--R
--R
--R   (2)  f
--R                                                          Type: BasicOperator
--E 30
--S 31 of 177
g := operator 'g
--R
--R
--R   (3)  g
--R                                                          Type: BasicOperator
--E 31
--S 32 of 177
ode := f(x*D(y x,x) - y(x)) = g(D(y x,x))
--R
--R
--R            ,                ,
--R   (4)  f(xy (x) - y(x))= g(y (x))
--R
--R                                          Type: Equation(Expression(Integer))
--E 32
--S 33 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseODE: equation has order 0
--R
--R   Continuing to read the file...
--R
--E 33
)clear all

--S 34 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 34
--S 35 of 177
ode := D(y x,x)  = (3*x^2-y(x)^2-7)/(exp(y(x))+2*x*y(x)+1)
--R
--R
--R                       2     2
--R         ,       - y(x)  + 3x  - 7
--R   (2)  y (x)= --------------------
--R                 y(x)
--R               %e     + 2x y(x) + 1
--R                                          Type: Equation(Expression(Integer))
--E 35
--S 36 of 177
solve(ode,y,x)
--R
--R
--R          y(x)         2           3
--R   (3)  %e     + x y(x)  + y(x) - x  + 7x
--R                                         Type: Union(Expression(Integer),...)
--E 36
)clear all

--S 37 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 37
--S 38 of 177
ode := D(y x,x)  = (2*x^3*y(x) - (y x)^4)/(x^4 - 2*x*(y x)^3)
--R
--R
--R                   4     3
--R         ,     y(x)  - 2x y(x)
--R   (2)  y (x)= ---------------
--R                       3    4
--R                2x y(x)  - x

--R                                          Type: Equation(Expression(Integer))
--E 38
--S 39 of 177
solve(ode,y,x)
--R
--R
--R   (3)  "failed"
--R                                                    Type: Union("failed",...)
--E 39
)clear all

--S 40 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 40
--S 41 of 177
ode := D(y x,x)*(D(y x,x) + y(x)) = x*(x + y(x))
--R
--R
--R         ,   2        ,               2
--R   (2)  y (x)  + y(x)y (x)= x y(x) + x
--R
--R                                          Type: Equation(Expression(Integer))
--E 41
--S 42 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 42
)clear all

--S 43 of 177
y := operator 'y
--R

--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 43
--S 44 of 177
ode := D(y x,x) = x/(x^2*(y x)^2 + (y x)^5)
--R
--R
--R         ,            x
--R   (2)  y (x)= ---------------
--R                   5    2    2
--R               y(x)  + x y(x)
--R                                          Type: Equation(Expression(Integer))
--E 44
--S 45 of 177
solve(ode,y,x)
--R
--R
--R                                     3
--R                                2y(x)
--R                              - ------
--R                3     2            3
--R        (- 2y(x)  - 2x  - 3)%e
--R   (3)  ------------------------------
--R                       4
--R                                         Type: Union(Expression(Integer),...)
--E 45
)clear all

--S 46 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 46
--S 47 of 177
ode := y(x) = 2*x*D(y x,x) - a*D(y x,x)^3
--R
--R
--R                   ,   3      ,
--R   (2)  y(x)= - a y (x)  + 2xy (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 47

--S 48 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 48
)clear all

--S 49 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 49
--S 50 of 177
ode := y(x) = 2*x*D(y x,x) - D(y x,x)^3
--R
--R
--R                 ,   3      ,
--R   (2)  y(x)= - y (x)  + 2xy (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 50
--S 51 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 51
)clear all

--S 52 of 177

y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 52
--S 53 of 177
ode := D(y x,x) = exp(x)*(y x)^2 - y(x) + exp(-x)
--R
--R
--R         ,         2  x     - x
--R   (2)  y (x)= y(x) %e  + %e    - y(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 53
--S 54 of 177
solve(ode,y,x)
--R
--R
--R   (3)  "failed"
--R                                                    Type: Union("failed",...)
--E 54
)clear all

--S 55 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 55
--S 56 of 177
ode := D(y x,x) = (y x)^2 - x*y(x) + 1
--R
--R
--R         ,         2
--R   (2)  y (x)= y(x)  - x y(x) + 1
--R
--R                                          Type: Equation(Expression(Integer))
--E 56
--S 57 of 177
solve(ode,y,x)
--R
--R

--R                         2
--R                        x
--R                      - --   x
--R                         2 ++       1
--I        (- y(x) + x)%e     |   - ------- d%J  + 1
--R                          ++           2
--I                                     %J
--R                                   - ---
--R                                      2
--R                                 %e
--R   (3)  -----------------------------------------
--R                                    2
--R                                   x
--R                                 - --
--R                                    2
--R                     (y(x) - x)%e
--R                                         Type: Union(Expression(Integer),...)
--E 57
)clear all

--S 58 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 58
--S 59 of 177
ode := D(y x,x) = (9*x^8 + 1)/((y x)^2 +1)
--R
--R
--R                  8
--R         ,      9x  + 1
--R   (2)  y (x)= ---------
--R                   2
--R               y(x)  + 1
--R                                          Type: Equation(Expression(Integer))
--E 59
--S 60 of 177
solve(ode,y,x)
--R
--R
--R            3             9
--R        y(x)  + 3y(x) - 3x  - 3x
--R   (3)  ------------------------
--R                    3

--R                                         Type: Union(Expression(Integer),...)
--E 60
)clear all

--S 61 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 61
--S 62 of 177
ode := y(x)=2*x*D(y x,x) + y(x)*D(y x,x)^2
--R
--R
--R                   ,   2      ,
--R   (2)  y(x)= y(x)y (x)  + 2xy (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 62
--S 63 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 63
)clear all

--S 64 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 64
--S 65 of 177
ode := x = y(x)*D(y x,x) - x*D(y x,x)^2
--R

--R
--R                ,   2        ,
--R   (2)  x= - x y (x)  + y(x)y (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 65
--S 66 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 66
-- Second Order Equations
)clear all
--S 67 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 67
--S 68 of 177
ode := D(y x,x,2)*(a*x+b)^2+4*D(y x,x)*(a*x+b)*a+2*y(x)*a^2=0
--R
--R
--R          2 2             2  ,,         2          ,        2
--R   (2)  (a x  + 2a b x + b )y  (x) + (4a x + 4a b)y (x) + 2a y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 68
--S 69 of 177
solve(ode,y,x)
--R
--R
--R                                        x              2a x + b
--R   (3)  [particular= 0,basis= [------------------,------------------]]
--R                                2 2             2  2 2             2
--R                               a x  + 2a b x + b  a x  + 2a b x + b
--IType: Union(Record(particular: Expression(Integer),...

--E 69
)clear all
--S 70 of 177
u := operator 'u
--R
--R
--R   (1)  u
--R                                                          Type: BasicOperator
--E 70
--S 71 of 177
ode := (x^2 - x)*D(u x,x,2) + (2*x^2+4*x-3)*D(u x,x) + 8*x*u(x)=1
--R
--R
--R          2      ,,         2           ,
--R   (2)  (x  - x)u  (x) + (2x  + 4x - 3)u (x) + 8x u(x)= 1
--R
--R                                          Type: Equation(Expression(Integer))
--E 71
--S 72 of 177
solve(ode,u,x)
--R
--R
--R                         3     2                                   - 2x
--R                       2x  - 3x  + 53                 1          %e
--R   (3)  [particular= ------------------,basis= [-------------,-----------]]
--R                        4      3      2          4     3    2  2
--R                     12x  - 24x  + 12x          x  - 2x  + x  x  - 2x + 1
--IType: Union(Record(particular: Expression(Integer),...
--E 72
)clear all
--S 73 of 177
w := operator 'w
--R
--R
--R   (1)  w
--R                                                          Type: BasicOperator
--E 73
--S 74 of 177
ode := (x^2 - x)*D(w x,x,2) + (1-2*x^2)*D(w x,x) + (4*x - 2)*w(x) = 0
--R
--R
--R          2      ,,           2      ,
--R   (2)  (x  - x)w  (x) + (- 2x  + 1)w (x) + (4x - 2)w(x)= 0

--R
--R                                          Type: Equation(Expression(Integer))
--E 74
--S 75 of 177
solve(ode,w,x)
--R
--R
--R                                2   2x
--R   (3)  [particular= 0,basis= [x ,%e  ]]
--IType: Union(Record(particular: Expression(Integer),...
--E 75
)clear all
--S 76 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 76
--S 77 of 177
ode := D(y x,x,2) - D(y x,x) = 2*y(x)*D(y x,x)
--R
--R
--R         ,,       ,           ,
--R   (2)  y  (x) - y (x)= 2y(x)y (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 77
--S 78 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getfreelincoeff: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 78
)clear all
--S 79 of 177
y := operator 'y
--R

--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 79
--S 80 of 177
ode := D(y x,x,2)/y(x) - D(y x,x)^2/y(x)^2 -1 + y(x)^(-3) = 0
--R
--R
--R            2 ,,           ,   2       3
--R        y(x) y  (x) - y(x)y (x)  - y(x)  + 1
--R
--R   (2)  ------------------------------------= 0
--R                            3
--R                        y(x)
--R                                          Type: Equation(Expression(Integer))
--E 80
--S 81 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 81
)clear all

--S 82 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 82
--S 83 of 177
ode :=  D(y x,x,2) + 2*x*D(y x,x) = 2*x
--R
--R
--R         ,,         ,
--R   (2)  y  (x) + 2xy (x)= 2x
--R
--R                                          Type: Equation(Expression(Integer))
--E 83

--S 84 of 177
solve(ode,y,x)
--R
--R
--R   (3)  [particular= x,basis= [1,erf(x)]]
--IType: Union(Record(particular: Expression(Integer),...
--E 84
)clear all
--S 85 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 85
--S 86 of 177
ode := 2*y(x)*D(y x,x,2) - D(y x,x)^2 = (D(y x,x) - x*D(y x,x,2))^2/3
--R
--R
--R                               2 ,,   2      ,    ,,       ,   2
--R                              x y  (x)  - 2xy (x)y  (x) + y (x)
--R              ,,       ,   2
--R   (2)  2y(x)y  (x) - y (x) = ----------------------------------
--R                                               3
--R                                          Type: Equation(Expression(Integer))
--E 86
--S 87 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 87
)clear all

--S 88 of 177
y := operator 'y
--R
--R

--R   (1)  y
--R                                                          Type: BasicOperator
--E 88
--S 89 of 177
ode := x*D(y x,x,2) = 2*y(x)*D(y x,x)
--R
--R
--R          ,,           ,
--R   (2)  xy  (x)= 2y(x)y (x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 89
--S 90 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getfreelincoeff: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 90
)clear all

--S 91 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 91
--S 92 of 177
ode := (1-x)*(y(x)*D(y x,x,2) - D(y x,x)^2) + x^2*y(x)^2 = 0
--R
--R
--R                      ,,              ,   2    2    2
--R   (2)  (- x + 1)y(x)y  (x) + (x - 1)y (x)  + x y(x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 92
--S 93 of 177
solve(ode,y,x)
--R

--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 93
)clear all
--S 94 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 94
--S 95 of 177
ode := x*y(x)*D(y x,x,2) + x*D(y x,x)^2 + y(x)*D(y x,x) = 0
--R
--R
--R               ,,         ,   2        ,
--R   (2)  x y(x)y  (x) + x y (x)  + y(x)y (x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 95
--S 96 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 96
)clear all
--S 97 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 97

--S 98 of 177
ode := D(y x,x,2)^2 - 2*D(y x,x,2)*D(y x,x) + 2*y(x)*D(y x,x) -y(x)^2 = 0
--R
--R
--R         ,,   2     ,    ,,            ,          2
--R   (2)  y  (x)  - 2y (x)y  (x) + 2y(x)y (x) - y(x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 98
--S 99 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 99
)clear all
--S 100 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 100
--S 101 of 177
ode := (x^3/2-x^2)*D(y x,x,2) + (2*x^2-3*x+1)*D(y x,x) + (x-1)*y(x) = 0
--R
--R
--R          3     2  ,,         2           ,
--R        (x  - 2x )y  (x) + (4x  - 6x + 2)y (x) + (2x - 2)y(x)
--R
--R   (2)  -----------------------------------------------------= 0
--R                                  2
--R                                          Type: Equation(Expression(Integer))
--E 101
--S 102 of 177
solve(ode,y,x)
--R
--R
--R   (3)

--R   [particular= 0,
--R                1               1                  1
--R              - - +-------+   - - +-------+   x   -- +----------+
--I                x |   1         x |   1     ++    %J |     1
--I    basis= [%e    |------- ,%e    |-------  |   %e   |---------- d%J ]]
--R                  | 2             | 2      ++        |  4      3
--I                 \|x  - 2x       \|x  - 2x          \|%J  - 2%J
--IType: Union(Record(particular: Expression(Integer),...
--E 102
)clear all
--S 103 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 103
--S 104 of 177
ode := D(y x,x,2) - 2*x*D(y x,x) + 2*y(x) = 3
--R
--R
--R         ,,         ,
--R   (2)  y  (x) - 2xy (x) + 2y(x)= 3
--R
--R                                          Type: Equation(Expression(Integer))
--E 104
--S 105 of 177
solve(ode,y,x)
--R
--R
--R                                           2
--I                                     x   %J
--R                     3             ++  %e
--I   (3)  [particular= -,basis= [x,x |   ----- d%J ]]
--R                     2            ++      2
--I                                        %J
--IType: Union(Record(particular: Expression(Integer),...
--E 105
)clear all
--S 106 of 177
y := operator 'y
--R
--R
--R   (1)  y

--S 107 of 177
ode := sqrt(x)*D(y x,x,2) + 2*x*D(y x,x) + 3*y(x) = 0
--R
--R
--R         +-+ ,,         ,
--R   (2)  \|x y  (x) + 2xy (x) + 3y(x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 107
--S 108 of 177
solve(ode,y,x)
--R
--R
--R   (3)  [particular= 0,basis= []]
--IType: Union(Record(particular: Expression(Integer),...
--E 108
)clear all
--S 109 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 109
--S 110 of 177
ode := x^2*D(y x,x,2) + 3*x*D(y x,x) = 1/(x^4*y(x)^3)
--R
--R
--R         2 ,,         ,        1
--R   (2)  x y  (x) + 3xy (x)= -------
--R                             4    3
--R                            x y(x)
--R                                          Type: Equation(Expression(Integer))
--E 110
--S 111 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R

--R   Continuing to read the file...
--R
--E 111
)clear all
--S 112 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 112
--S 113 of 177
ode := D(y x,x,2) - 2/x^2*y(x) = 7*x^4 +3*x^3
--R
--R
--R         2 ,,
--R        x y  (x) - 2y(x)
--R                            4     3
--R   (2)  ----------------= 7x  + 3x
--R                2
--R               x
--R                                          Type: Equation(Expression(Integer))
--E 113
--S 114 of 177
solve(ode,y,x)
--R
--R
--R                       7     6     3               3      3
--R                     3x  + 2x  - 7x  + 14         x  - 1 x  + 2
--R   (3)  [particular= --------------------,basis= [------,------]]
--R                              12x                    x      x
--IType: Union(Record(particular: Expression(Integer),...
--E 114
)clear all
--S 115 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 115
--S 116 of 177
ode := D(y x,x,2) +y(x) = csc(x)

--R
--R
--R         ,,
--R   (2)  y  (x) + y(x)= csc(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 116
--S 117 of 177
solve(ode,y,x)
--R
--R
--R   (3)
--R                            sin(x)                     2
--R   [particular= sin(x)log(----------) - sin(x)log(----------) - x cos(x),
--R                          cos(x) + 1              cos(x) + 1
--R    basis= [cos(x),sin(x)]]
--IType: Union(Record(particular: Expression(Integer),...
--E 117
-- Higher Order Equations
)clear all
--S 118 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 118
--S 119 of 177
ode := D(y x,x,7) - 14*D(y x,x,6) +80*D(y x,x,5) -242*D(y x,x,4) + _
         419*D(y x,x,3) - 416*D(y x,x,2) +220*D(y x,x) -48*y(x) = 0
--R
--R
--R   (2)
--R      (vii)         (vi)         (v)          (iv)          ,,,          ,,
--R     y     (x) - 14y    (x) + 80y   (x) - 242y    (x) + 419y   (x) - 416y  (x)
--R
--R   +
--R         ,
--R     220y (x) - 48y(x)
--R
--R     =
--R     0
--R                                          Type: Equation(Expression(Integer))
--E 119

--S 120 of 177
solve(ode,y,x)
--R
--R
--R                                 4x   3x   2x     2x   x     x  2  x
--R   (3)  [particular= 0,basis= [%e  ,%e  ,%e  ,x %e  ,%e ,x %e ,x %e ]]
--IType: Union(Record(particular: Expression(Integer),...
--E 120
)clear all
--S 121 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 121
--S 122 of 177
ode := D(y x,x,4) -4/x^2*D(y x,x,2) + 8/x^3*D(y x,x) -8/x^4*D(y x,x) = 0
--R
--R
--R         4 (iv)        2 ,,               ,
--R        x y    (x) - 4x y  (x) + (8x - 8)y (x)
--R
--R   (2)  --------------------------------------= 0
--R                           4
--R                          x
--R                                          Type: Equation(Expression(Integer))
--E 122
--S 123 of 177
solve(ode,y,x)
--R
--R
--R   (3)  [particular= 0,basis= [1]]
--IType: Union(Record(particular: Expression(Integer),...
--E 123
)clear all
--S 124 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 124

--S 125 of 177
ode := (1+x+x^2)*D(y x,x,3) + (3+6*x)*D(y x,x,2) +6*D(y x,x) = 6*x
--R
--R
--R          2          ,,,               ,,        ,
--R   (2)  (x  + x + 1)y   (x) + (6x + 3)y  (x) + 6y (x)= 6x
--R
--R                                          Type: Equation(Expression(Integer))
--E 125
--S 126 of 177
solve(ode,y,x)
--R
--R
--R                         4
--R                        x  - 4                 1        x + 1
--R   (3)  [particular= ------------,basis= [----------,----------,1]]
--R                       2                   2          2
--R                     4x  + 4x + 4         x  + x + 1 x  + x + 1
--IType: Union(Record(particular: Expression(Integer),...
--E 126
)clear all
--S 127 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 127
--S 128 of 177
ode := (D(y x,x)^2 +1)*D(y x,x,3) - 3*D(y x,x)*D(y x,x,2) = 0
--R
--R
--R          ,   2      ,,,        ,    ,,
--R   (2)  (y (x)  + 1)y   (x) - 3y (x)y  (x)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 128
--S 129 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R

)clear all
--S 130 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 130
--S 131 of 177
ode := 3*D(y x,x,2)*D(y x,x,4) - 5*D(y x,x,3)^2 = 0
--R
--R
--R          ,,    (iv)        ,,,   2
--R   (2)  3y  (x)y    (x) - 5y   (x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 131
--S 132 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseLODE: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 132
-- Special Equations
)clear all
--S 133 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 133
--S 134 of 177
ode := D(y t,t) + a*y(t-1) = 0

--R
--R
--R         ,
--R   (2)  y (t) + a y(t - 1)= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 134
--S 135 of 177
solve(ode,y,t)
--R
--R
--R           t
--R         ++
--I   (3)   |   a y(%J - 1)d%J  + y(t)
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E 135
)clear all

--S 136 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 136
--S 137 of 177
ode := D(y(x,a),x) = a*y(x,a)
--R
--R
--R   (2)  y  (x,a)= a y(x,a)
--R         ,1
--R                                          Type: Equation(Expression(Integer))
--E 137
--S 138 of 177
solve(ode,y,x)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   parseODE: equation has order 0
--R
--R   Continuing to read the file...
--R
--E 138

)clear all

--S 139 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 139
--S 140 of 177
ode := D(y x,x,4) = sin(x)
--R
--R
--R         (iv)
--R   (2)  y    (x)= sin(x)
--R
--R                                          Type: Equation(Expression(Integer))
--E 140
--S 141 of 177
solve(ode,y,x=0,[0,0,0,0])
--R
--R
--R                   3
--R        6sin(x) + x  - 6x
--R   (3)  -----------------
--R                6
--R                                         Type: Union(Expression(Integer),...)
--E 141
)clear all
--S 142 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 142
--S 143 of 177
ode := x*D(y x,x,2) + D(y x,x) +2*x*y(x) =0
--R
--R
--R          ,,       ,
--R   (2)  xy  (x) + y (x) + 2x y(x)= 0
--R

--S 144 of 177
solve(ode,y,x=0,[1,0])
--R
--R
--R   (3)  "failed"
--R                                                    Type: Union("failed",...)
--E 144
)clear all
--S 145 of 177
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 145
--S 146 of 177
ode := x*D(y x,x)^2 -(y x)^2 + 1 = 0
--R
--R
--R           ,   2       2
--R   (2)  x y (x)  - y(x)  + 1= 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 146
--S 147 of 177
solve(ode,y,x=0,[1])
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 147
)clear all
--S 148 of 177
y := operator 'y
--R
--R
--R   (1)  y

--S 149 of 177
ode := D(y x,x,2) + y(x)*D(y x,x)^3 = 0
--R
--R
--R         ,,           ,   3
--R   (2)  y  (x) + y(x)y (x) = 0
--R
--R                                          Type: Equation(Expression(Integer))
--E 149
--S 150 of 177
solve(ode,y,x=0,[0,2])
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 150
-- Systems Of equations
)clear all
--S 151 of 177
x := operator 'x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 151
--S 152 of 177
y := operator 'y
--R
--R
--R   (2)  y
--R                                                          Type: BasicOperator
--E 152
--S 153 of 177
z := operator 'z
--R
--R
--R   (3)  z

--S 154 of 177
odes := [D(x t,t) = -3*y(t)*z(t), D(y t,t) = 3*x(t)*z(t), D(z t,t) = -x(t)*y(t)]
--R
--R
--R          ,                  ,                ,
--R   (4)  [x (t)= - 3y(t)z(t),y (t)= 3x(t)z(t),z (t)= - x(t)y(t)]
--R
--R                                    Type: List(Equation(Expression(Integer)))
--E 154
--S 155 of 177
solve(odes,[x,y,z],t)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getfreelincoeff: not a linear ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 155
)clear all
--S 156 of 177
x := operator 'x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 156
--S 157 of 177
y := operator 'y
--R
--R
--R   (2)  y
--R                                                          Type: BasicOperator
--E 157
--S 158 of 177
a := operator 'a
--R
--R
--R   (3)  a
--R                                                          Type: BasicOperator
--E 158

--S 159 of 177
b := operator 'b
--R
--R
--R   (4)  b
--R                                                          Type: BasicOperator
--E 159
--S 160 of 177
odes := [D(x t,t) = a(t)*((y t)^2 - (x t)^2) + 2*b(t)*x(t)*y(t) + 2*c*x(t),
         D(y t,t) = b(t)*((y t)^2 - (x t)^2) - 2*a(t)*x(t)*y(t) + 2*c*y(t)]
--R
--R
--R   (5)
--R     ,             2                           2
--R   [x (t)= a(t)y(t)  + 2b(t)x(t)y(t) - a(t)x(t)  + 2c x(t),
--R
--R     ,             2                                    2
--R    y (t)= b(t)y(t)  + (- 2a(t)x(t) + 2c)y(t) - b(t)x(t) ]
--R
--R                                    Type: List(Equation(Expression(Integer)))
--E 160
--S 161 of 177
solve(odes,[x,y],t)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   getlincoeff: not an appropriate ordinary differential equation
--R
--R   Continuing to read the file...
--R
--E 161
)clear all
--S 162 of 177
x := operator 'x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 162
--S 163 of 177
y := operator 'y
--R
--R

--S 164 of 177
odes := [D(x t,t) = x(t)*(1+cos(t)/(2+sin(t))), D(y t,t) = x(t) - y(t)]
--R
--R
--R          ,     x(t)sin(t) + x(t)cos(t) + 2x(t)  ,
--R   (3)  [x (t)= -------------------------------,y (t)= - y(t) + x(t)]
--R                           sin(t) + 2
--R                                    Type: List(Equation(Expression(Integer)))
--E 164
--S 165 of 177
solve(odes,[x,y],t)
--R
--R
--R   (4)  "failed"
--R                                                    Type: Union("failed",...)
--E 165
)clear all
--S 166 of 177
x := operator 'x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 166
--S 167 of 177
y := operator 'y
--R
--R
--R   (2)  y
--R                                                          Type: BasicOperator
--E 167
--S 168 of 177
odes := [D(x t,t) = 9*x(t) + 2*y(t), D(y t,t) = x(t) + 8*y(t)]
--R
--R
--R          ,                    ,
--R   (3)  [x (t)= 2y(t) + 9x(t),y (t)= 8y(t) + x(t)]
--R
--R                                    Type: List(Equation(Expression(Integer)))
--E 168

--S 169 of 177
solve(odes,[x,y],t)
--R
--R
--R                                            10t
--R                                      10t %e        7t     7t
--R   (4)  [particular= [0,0],basis= [[%e   ,-----],[%e  ,- %e  ]]]
--R                                            2
--IType: Union(Record(particular: Vector(Expression(Integer)),...
--E 169
)clear all
--S 170 of 177
x := operator 'x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 170
--S 171 of 177
y := operator 'y
--R
--R
--R   (2)  y
--R                                                          Type: BasicOperator
--E 171
--S 172 of 177
odes := [D(x t,t) - x(t) - 2*y(t) = 0, D(x t,t,2) - 2*D(y t,t) = 2*t - cos(2*t)]
--R
--R
--R          ,                       ,,        ,
--R   (3)  [x (t) - 2y(t) - x(t)= 0,x  (t) - 2y (t)= - cos(2t) + 2t]
--R
--R                                    Type: List(Equation(Expression(Integer)))
--E 172
--S 173 of 177
solve(odes,[x,y],t)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   solve: not a first order linear system
--R
--R   Continuing to read the file...
--R
--E 173

--S 174 of 177
y1 := operator 'y1
--R
--R
--R   (1)  y1
--R                                                          Type: BasicOperator
--E 174
--S 175 of 177
y2 := operator 'y2
--R
--R
--R   (2)  y2
--R                                                          Type: BasicOperator
--E 175
--S 176 of 177
odes := [D(y1 x,x) = -1/(x*(x^2 + 1))*y1(x) + 1/(x^2*(x^2 + 1))*y2(x)+1/x,_
         D(y2 x,x) = -x^2/(x^2 + 1)*y1(x) + (2*x^2+1)/x/(x^2+1)*y2(x)+1]
--R
--R
--R   (3)
--R                               3                2              3         3
--R      ,     y2(x) - x y1(x) + x  + x   ,     (2x  + 1)y2(x) - x y1(x) + x  + x
--R   [y1 (x)= ------------------------,y2 (x)= ---------------------------------]
--R                      4    2                                3
--R                     x  + x                                x  + x
--R                                    Type: List(Equation(Expression(Integer)))
--E 176
--S 177 of 177
solve(odes,[y1,y2],x)
--R
--R
--R                                                        1    2
--R   (4)  [particular= [log(x) - 1,x log(x) - x],basis= [[-,- x ],[1,x]]]
--R                                                        x
--IType: Union(Record(particular: Vector(Expression(Integer)),basis: ...
--E 177
)spool
)lisp (bye)

———-

References^{[1] nothing}

41